# Shrinking vs. expanding: the evolution of spatial support in degenerate Keller–Segel systems

### v1

@@ 1 Introduction
where $m>1$ is a given parameter and $\Omega\subset\mathbb{R}^{n}$ is a ball. The system (1.1) is a parabolic–elliptic simplification of the celebrated Keller–Segel model ([22]), which models the spatio-temporal evolution of an organism (with density $u$) that is partially attracted by a chemical (with concentration $v$) produced by itself. This chemotactic effect is modelled by the term $-\nabla\cdot(u\nabla v)$ and the non-directed mobility is supposed to be density-dependent and to take the form $\nabla\cdot(u^{m-1}\nabla u)$; see, e.g., [28], [17], [15] for modelling considerations.
@@
$A_{\mathrm{crit}}\coloneqq\left[\frac{\int_{\Omega}u_{0}\left(1-\frac{r_{1}^{n}}{R^{n}}\right)(m-1)}{\omega_{n}r_{1}^{n-1}n}\right]^{\frac{1}{m-1}}.$ (1.3)
@@
then the spatial support initially shrinks in the sense that there are $T>0$ and $\zeta>0$ with
@@
then the spatial support initially expands in the sense that there are $T>0$ and $\zeta>0$ with
@@
Since $(r_{1}-r_{0})^{\alpha-\frac{1}{m-1}}$ converges to 00 (respectively, $\infty$) as $r_{0}\nearrow r_{1}$ if $\alpha>\frac{1}{m-1}$ (respectively, $\alpha<\frac{1}{m-1}$), we see that
@@
    While the condition (1.9) with $\frac{1}{m-1}$ replaced by $\frac{2}{m-2}$ implies initial expanding of the support for solutions to the PME, the analogue of (1.8) does not imply initial support shrinking but instead that the support remains constant for some time. That is, the critical exponent for the PME distinguishes between the existence of waiting times and immediate expanding, the critical exponent for (1.1) between shrinking and expanding.
@@
    The behavior at the critical exponent is also different. While for (1.1), Theorem 1.1 shows the existence of a critical parameter $A_{\mathrm{crit}}$ and hence in particular that the support of solutions to initial data $u_{0}$ with $u_{0}(r)\sim(r_{1}-r)_{+}^{\frac{1}{m-1}}$ may shrink or expand depending on the implied constant, all solutions to the PME with initial data $u_{0}(r)\leq A(r_{1}-r)_{+}^{\frac{2}{m-1}}$ exhibit a waiting time phenomenon, regardless of how large $A$ is ([1, Propsition 4.2]).
@@
For the variant of (1.1) posed in the full space and with the second equation replaced by $-\Delta v=u$, it has been claimed (without detailed proof) in [14, Corollary on p. 1613] that (1.8) implies initial shrinking, provided that $m>2-\frac{2}{n}$. We again note a few difference to the situation in Theorem 1.1.
@@
$\partial_{t}w-n^{2}s^{2-\frac{2}{n}}(\partial_{s}w)^{m-1}\partial_{ss}w-w\partial_{s}w+\mu s\partial_{s}w=0\qquad\text{in $(0,R^{n})\times(0,T)$}$ (1.10)
@@
$\big{(}\sup\{\,r\in(0,R)\mid u(r,t)>0\,\}\big{)}^{n}=\inf\{\,s\in(0,R^{n})\mid w(s,t)=\mu R^{n}\,\}\eqqcolon I_{w}(t)\qquad\text{for all $t\in(0,T)$},$ (1.11)
@@ 2 Local existence and convergence to weak solutions
Let $T>0$ and let $u_{0}\in L^{\infty}(\Omega)$ be nonnegative and radially symmetric with $\frac{1}{|\Omega|}\int_{\Omega}u_{0}=\mu$. We call a pair of radially symmetric functions
$(u,v)\in L_{\mathrm{loc}}^{\infty}(\overline{\Omega}\times[0,T))\times L_{\mathrm{loc}}^{\infty}([0,T);W^{1,2}(\Omega))\quad\text{with}\quad u^{m}\in L_{\mathrm{loc}}^{2}([0,T);W^{1,2}(\Omega))$
being such that $u\geq 0$ a.e. in $\Omega\times(0,T)$ a radial weak solution of (1.1) if
@@
hold for all $\varphi\in C_{c}^{\infty}(\overline{\Omega}\times[0,T))$.
@@
Let $T>0$, let $u_{0}\in L^{\infty}(\Omega)$ be nonnegative and radially symmetric and let $(u,v)$ be a radial weak solution of (1.1) in the sense of Definition 2.1.
@@
    As also noted in [30, Remark after Definition 2.1], the weak formulation (2.1) implies that $u_{t}\in L_{\mathrm{loc}}^{2}([0,T);(W^{1,2}(\Omega))^{\star})$ and hence (after redefining $u$ on a null set of times) $u\in C^{0}([0,T);L^{2}(\Omega))$.
@@
    It has recently been shown in [5, Corollary 1.9] that $u$ is locally (space-time) Hölder continuous in $\overline{\Omega}\times(0,T)$ and that Hölder continuity of $u_{0}$ implies that $u$ is Hölder continuous up to $t=0$. (For $m\leq 3$, the same conclusion can be obtained from the classical work [29].)
@@
Let $u_{0}\in L^{\infty}(\Omega)$ be nonnegative and radially symmetric. Then we can find $T>0$ such that for all $\varepsilon\in(0,1)$, a classical solution $(u_{\varepsilon},v_{\varepsilon})$ of
$\begin{cases}u_{\varepsilon t}=\nabla\cdot((u_{\varepsilon}+\varepsilon)^{m-1}\nabla u_{\varepsilon}-u_{\varepsilon}\nabla v_{\varepsilon})&\text{in $\Omega\times(0,T)$},\\ 0=\Delta v_{\varepsilon}-\mu+u_{\varepsilon},\quad\mu=\frac{1}{|\Omega|}\int_{\Omega}u_{0},&\text{in $\Omega\times(0,T)$},\\ \partial_{\nu}u_{\varepsilon}=\partial_{\nu}v_{\varepsilon}=0&\text{on $\partial\Omega\times(0,T)$},\\ u_{\varepsilon}(\cdot,0)=u_{0}&\text{in $\Omega$}\end{cases}$ (2.2)
@@
$u_{\varepsilon}\in C^{2,1}(\overline{\Omega}\times(0,T))\cap C^{0}([0,T);L^{1}(\Omega)),\quad v_{\varepsilon}\in C^{2,0}(\overline{\Omega}\times(0,T))\quad\text{and}\quad\int_{\Omega}v_{\varepsilon}=0$
exists, which is moreover such that $u_{\varepsilon}(\cdot,t)$ and $v_{\varepsilon}(\cdot,t)$ are radially symmetric for all $t\in(0,T)$ and that
$0<u_{\varepsilon}\leq\|u_{0}\|_{L^{\infty}(\Omega)}+1\qquad\text{in $\overline{\Omega}\times(0,T)$}.$ (2.3)
@@
Let $u_{0}\in L^{\infty}(\Omega)$ be nonnegative and radially symmetric and let $T$ and $(u_{\varepsilon},v_{\varepsilon})_{\varepsilon\in(0,1)}$ be as given by Lemma 2.3. Then there exist a null sequence $(\varepsilon_{j})_{j\in\mathbb{N}}\subset(0,1)$ and a radial weak solution $(u,v)$ of (1.1) in the sense of Definition 2.1 such that
[…]
  • Proof.
@@
Let $u_{0}\in L^{\infty}(\Omega)$ be nonnegative and radially symmetric and let $T$ and $(u_{\varepsilon},v_{\varepsilon})_{\varepsilon\in(0,1)}$ be as given by Lemma 2.3. Then for every $r\in(0,R)$ and $\delta>0$, there exists $t_{0}\in(0,T)$ such that
@@
    for all $\varphi\in C^{\infty}_{c}(\Omega)$, $t\in(0,T)$ and $\varepsilon\in(0,1)$. By for instance testing the second equation in (2.2) with $v_{\varepsilon}$ to gain a uniform-in-time gradient bound for $v_{\varepsilon}$ combined with the uniform bound for $u_{\varepsilon}$ found in (2.3), we gain $C\equiv C(\|u_{0}\|_{L^{\infty}(\Omega)})>0$ such that
@@
    for all $\varphi\in C^{\infty}_{c}(\Omega)$, $t\in(0,T)$ and $\varepsilon\in(0,1)$.
@@
For the remainder of the paper, we now fix a nonnegative, radially symmetric $u_{0}\in L^{\infty}(\Omega)$ and set $\mu\coloneqq\frac{1}{|\Omega|}\int_{\Omega}u_{0}$. Moreover, we fix the family of approximate solutions $(u_{\varepsilon},v_{\varepsilon})_{\varepsilon\in(0,1)}$ constructed in Lemma 2.3, denote the uniform time given by Lemma 2.3 by $T_{0}>0$ and fix the limit solution $(u,v)$ found in Lemma 2.4.
@@ 3 Mass accumulation functions
on $[0,R^{n}]\times[0,T_{0}]$ with initial data $w_{0}(s)\coloneqq n\int_{0}^{s^{\frac{1}{n}}}\rho^{n-1}u_{0}(\rho)\,\mathrm{d}\rho$. Additionally, it is immediately obvious from the nonnegativity of $u_{\varepsilon}$ as well as the mass conservation properties of (2.2) that $w_{\varepsilon}(\cdot,t)$ is monotonically increasing and $w_{\varepsilon}(0,t)=0$ as well as $w_{\varepsilon}(R^{n},t)=\mu R^{n}$ for all $t\in[0,T_{0}]$.
@@ 4 Comparison argument
$\underline{w}(s,0)-\overline{w}(s,0)$ $\leq 0$ for all $\qquad\text{for all $s\in(a,b)$},$ (4.7)
$\underline{w}(s,t)-\overline{w}(s,t)$ $\leq 0$ for all s{a,b} and $\qquad\text{for all $s\in\{a,b\}$ and $t\in(0,T)$},$ (4.8)
then
$\underline{w}(s,t)-\overline{w}(s,t)$ $\leq 0$ for all s(a,b) and $\qquad\text{for all $s\in(a,b)$ and $t\in(0,T)$}.$ (4.9)
  • Proof.
    $\overline{W}(s,t)\coloneqq\overline{w}\left(\tfrac{b-a}{b}s+a,t\right)+\varepsilon\tfrac{b-a}{b}s\quad\text{ and }\quad\underline{W}(s,t)\coloneqq\underline{w}\left(\tfrac{b-a}{b}s+a,t\right)+\varepsilon\tfrac{b-a}{b}s$
    for all $(s,t)\in[0,L]\times[0,T)$ with $L\coloneqq b$. As we have only modified the original functions $\overline{w}$, $\underline{w}$ in a linear fashion, the new functions $\overline{W}$, $\underline{W}$ naturally retain all of our assumed regularity properties as well as the proper ordering at the parabolic boundary of $[0,L]\times[0,T)$ due to (4.7) as well as (4.8). Due to (4.3) and $a<b$, it is further directly evident that $\partial_{s}\overline{W}>0$ and $\partial_{s}\underline{W}>0$ in $(0,L)\times(0,T)$. Finally due to (4.5) and (4.6), it follows by direct computation that
    $\partial_{t}\underline{W}\leq\Phi(s,t,\underline{W},\partial_{s}\underline{W},\partial_{ss}\underline{W})\quad\text{ and }\quad\partial_{t}\overline{W}\geq\Phi(s,t,\overline{W},\partial_{s}\overline{W},\partial_{ss}\overline{W})$
@@ 4.1 Case of support shrinking
    for all $r_{0}\in[r_{\mathrm{min}},r_{1})$, and such that the quantities $\eta$ and $\delta$ defined in (4.10) fulfill
@@
    for all $t\in[0,\infty)$. Moreover, $\partial_{ss}\underline{w}_{\mathrm{mid}}$ is continuous on the compact set $[r_{0}^{n},\rho(t)-\delta]$ for all $t\in[0,\infty)$, so that all our desired regularity properties for $\underline{w}$ are evidently fulfilled.
@@
    $-\underline{w}_{\mathrm{mid}}+\mu s\leq-\theta_{\mathrm{max}}-\varepsilon_{0}\kappa(R^{n}-r_{1}^{n})-(1-\lambda)\mu(R^{n}-r_{1}^{n})$ (4.20)
@@
    For our first comparison function, let now $\underline{w}_{\varepsilon,\mathrm{stat}}$ be as in (4.11) for all $\varepsilon\in(0,\varepsilon_{1})$ with the parameters as fixed above but $\theta$ replaced by 00. Then we gain
@@
    Again due to $\eta\geq 0$, $\lambda>1$ as well as the fact that $w_{0}(s)\geq\mu R^{n}-C(r_{1}^{n}-s)_{+}^{\frac{m}{m-1}}\eqqcolon z_{0}(s)$ for $s\in[r_{0}^{n},R^{n}]$ by (4.21), we see that $w_{0}(s)\geq\underline{w}_{\varepsilon,\mathrm{stat}}(s,0)$ for all $s\in[r_{0}^{n},r_{1}^{n}-\delta]\cup[r_{1}^{n},R^{n}]$. Moreover, since $w_{0}\geq z_{0}$ and $z_{0}(r_{1}^{n})\geq\underline{w}_{\varepsilon,\mathrm{stat}}(r_{1}^{n},0)$, the ordering $w_{0}(s)\geq\underline{w}_{\varepsilon,\mathrm{stat}}(s,0)$ in $[r_{1}^{n}-\delta,r_{1}^{n}]$ and hence on the full interval $[r_{0}^{n},R^{n}]$ follows if $\underline{w}_{\varepsilon,\mathrm{stat}}(\cdot,0)-z_{0}$ is increasing in the former interval. Indeed,
@@
    Notably when $\theta=0$, the condition (4.12) is always trivially fulfilled. Thus Lemma 4.2 ensures that we can apply the comparison result in Lemma 4.1 to $w_{\varepsilon}$ and $\underline{w}_{\varepsilon,\mathrm{stat}}$ and thus gain $w_{\varepsilon}\geq\underline{w}_{\varepsilon,\mathrm{stat}}$ on $[r_{0}^{n},R^{n}]\times[0,T]$. This then directly yields
@@
    a.e. on $[r_{0}^{n},r_{2}^{n}]\times[0,T]$ due to Lemma 4.2 and our previous parameter choices. As initial data ordering follows by the same argument as before, this again allows us to apply Lemma 4.1 to gain that
@@
    on $[r_{0}^{n},r_{1}^{n}]\times[0,T]$. As $\underline{w}_{\varepsilon,\mathrm{shrink}}$ converges pointwise to

### v2

@@ 1 Introduction
where $m>1$ is a given parameter and $\Omega\subset\mathbb{R}^{n}$ is a ball. The system (1.1) is a parabolic–elliptic simplification of the celebrated Keller–Segel model ([22]), which models the spatio-temporal evolution of an organism (with density $u$) that is partially attracted by a chemical (with concentration $v$) produced by itself. This chemotactic effect is modelled by the term $\nabla\cdot(u\nabla v)$ and the non-directed mobility is supposed to be density-dependent and to take the form $\nabla\cdot(u^{m-1}\nabla u)$; see, e.g., [28], [17], [15] for modelling considerations.
@@
$A_{\mathrm{crit}}\coloneqq\left[\frac{\int_{\Omega}u_{0}\left(1-\frac{r_{1}^{n}}{R^{n}}\right)(m-1)}{\omega_{n}r_{1}^{n-1}}\right]^{\frac{1}{m-1}}.$ (1.3)
@@
then the spatial support initially shrinks in the sense that there are $T\in(0,T_{0})$ and $\zeta>0$ with
@@
then the spatial support initially expands in the sense that there are $T\in(0,T_{0})$ and $\zeta>0$ with
@@
Since $(r_{1}-r_{0})^{\alpha-\frac{1}{m-1}}$ converges to 0 (respectively, $\infty$) as $r_{0}\nearrow r_{1}$ if $\alpha>\frac{1}{m-1}$ (respectively, $\alpha<\frac{1}{m-1}$), we see that
@@
    While the condition (1.9) with $\frac{1}{m-1}$ replaced by $\frac{2}{m-1}$ implies initial expanding of the support for solutions to the PME, the analogue of (1.8) does not imply initial support shrinking but instead that the support remains constant for some time. That is, the critical exponent for the PME distinguishes between the existence of waiting times and immediate expanding, the critical exponent for (1.1) between shrinking and expanding.
@@
    The behaviour at the critical exponent is also different. While for (1.1), Theorem 1.1 shows the existence of a critical parameter $A_{\mathrm{crit}}$ and hence in particular that the support of solutions to initial data $u_{0}$ with $u_{0}(r)\sim(r_{1}-r)_{+}^{\frac{1}{m-1}}$ may shrink or expand depending on the implied constant, all solutions to the PME with initial data $u_{0}(r)\leq A(r_{1}-r)_{+}^{\frac{2}{m-1}}$ exhibit a waiting time phenomenon, regardless of how large $A$ is ([1, Propsition 4.2]).
@@
For the variant of (1.1) posed in the full space and with the second equation replaced by $\Delta v=u$, it has been claimed (without detailed proof) in [14, Corollary on p. 1613] that (1.8) implies initial shrinking, provided that $m>2-\frac{2}{n}$. We again note a few difference to the situation in Theorem 1.1.
@@
$\partial_{t}w-n^{2}s^{2-\frac{2}{n}}(\partial_{s}w)^{m-1}\partial_{ss}w-w\partial_{s}w+\mu s\partial_{s}w=0\qquad\text{in $(0,R^{n})\times(0,T_{0})$}$ (1.10)
@@
$\big(\sup\{\,r\in(0,R)\mid u(r,t)>0\,\}\big)^{n}=\inf\{\,s\in(0,R^{n})\mid w(s,t)=\mu R^{n}\,\}\eqqcolon I_{w}(t)\qquad\text{for all $t\in(0,T_{0})$},$ (1.11)
@@ 2 Local existence and convergence to weak solutions
Let $T_{0}>0$ and let $u_{0}\in L^{\infty}(\Omega)$ be nonnegative and radially symmetric with $\frac{1}{|\Omega|}\int_{\Omega}u_{0}=\mu$. We call a pair of radially symmetric functions
$(u,v)\in L_{\mathrm{loc}}^{\infty}(\overline{\Omega}\times[0,T_{0}))\times L_{\mathrm{loc}}^{\infty}([0,T_{0});W^{1,2}(\Omega))\quad\text{with}\quad u^{m}\in L_{\mathrm{loc}}^{2}([0,T_{0});W^{1,2}(\Omega))$
being such that $u\geq 0$ a.e. in $\Omega\times(0,T_{0})$ a radial weak solution of (1.1) if
@@
hold for all $\varphi\in C_{c}^{\infty}(\overline{\Omega}\times[0,T_{0}))$.
@@
Let $T_{0}>0$, let $u_{0}\in L^{\infty}(\Omega)$ be nonnegative and radially symmetric and let $(u,v)$ be a radial weak solution of (1.1) in the sense of Definition 2.1.
@@
    As also noted in [30, Remark after Definition 2.1], the weak formulation (2.1) implies that $u_{t}\in L_{\mathrm{loc}}^{2}([0,T_{0});(W^{1,2}(\Omega))^{\star})$ and hence (after redefining $u$ on a null set of times) $u\in C^{0}([0,T_{0});L^{2}(\Omega))$.
@@
    It has recently been shown in [5, Corollary 1.9] that $u$ is locally (space-time) Hölder continuous in $\overline{\Omega}\times(0,T_{0})$ and that Hölder continuity of $u_{0}$ implies that $u$ is Hölder continuous up to $t=0$. (For $m\leq 3$, the same conclusion can be obtained from the classical work [29].)
@@
Let $u_{0}\in L^{\infty}(\Omega)$ be nonnegative and radially symmetric. Then we can find $T_{0}>0$ such that for all $\varepsilon\in(0,1)$, a classical solution $(u_{\varepsilon},v_{\varepsilon})$ of
$\begin{cases}u_{\varepsilon t}=\nabla\cdot((u_{\varepsilon}+\varepsilon)^{m-1}\nabla u_{\varepsilon}-u_{\varepsilon}\nabla v_{\varepsilon})&\text{in $\Omega\times(0,T_{0})$},\\ 0=\Delta v_{\varepsilon}-\mu+u_{\varepsilon},\quad\mu=\frac{1}{|\Omega|}\int_{\Omega}u_{0},&\text{in $\Omega\times(0,T_{0})$},\\ \partial_{\nu}u_{\varepsilon}=\partial_{\nu}v_{\varepsilon}=0&\text{on $\partial\Omega\times(0,T_{0})$},\\ u_{\varepsilon}(\cdot,0)=u_{0}&\text{in $\Omega$}\end{cases}$ (2.2)
@@
$u_{\varepsilon}\in C^{2,1}(\overline{\Omega}\times(0,T_{0}))\cap C^{0}([0,T_{0});L^{1}(\Omega)),\quad v_{\varepsilon}\in C^{2,0}(\overline{\Omega}\times(0,T_{0}))\quad\text{and}\quad\int_{\Omega}v_{\varepsilon}=0$
exists, which is moreover such that $u_{\varepsilon}(\cdot,t)$ and $v_{\varepsilon}(\cdot,t)$ are radially symmetric for all $t\in(0,T_{0})$ and that
$0<u_{\varepsilon}\leq\|u_{0}\|_{L^{\infty}(\Omega)}+1\qquad\text{in $\overline{\Omega}\times(0,T_{0})$}.$ (2.3)
@@
Let $u_{0}\in L^{\infty}(\Omega)$ be nonnegative and radially symmetric and let $T_{0}$ and $(u_{\varepsilon},v_{\varepsilon})_{\varepsilon\in(0,1)}$ be as given by Lemma 2.3. Then there exist a null sequence $(\varepsilon_{j})_{j\in\mathbb{N}}\subset(0,1)$ and a radial weak solution $(u,v)$ of (1.1) in the sense of Definition 2.1 such that
$u_{\varepsilon}$ $\to u$ $\qquad\text{a.e.\ in }\Omega\times(0,T_{0})\text{ as $\varepsilon=\varepsilon_{j}\searrow 0$}.$ (2.4)
@@
Let $u_{0}\in L^{\infty}(\Omega)$ be nonnegative and radially symmetric and let $T_{0}$ and $(u_{\varepsilon},v_{\varepsilon})_{\varepsilon\in(0,1)}$ be as given by Lemma 2.3. Then for every $r\in(0,R)$ and $\delta>0$, there exists $t_{0}\in(0,T_{0})$ such that
@@
for all $\varphi\in C^{\infty}_{c}(\Omega)$, $t\in(0,T_{0})$ and $\varepsilon\in(0,1)$. By for instance testing the second equation in (2.2) with $v_{\varepsilon}$ to gain a uniform-in-time gradient bound for $v_{\varepsilon}$ combined with the uniform bound for $u_{\varepsilon}$ found in (2.3), we gain $C\equiv C(\|u_{0}\|_{L^{\infty}(\Omega)})>0$ such that
@@
for all $\varphi\in C^{\infty}_{c}(\Omega)$, $t\in(0,T_{0})$ and $\varepsilon\in(0,1)$.
@@
For the remainder of the paper, we now fix a nonnegative, radially symmetric $u_{0}\in L^{\infty}(\Omega)$ and set $\mu\coloneqq\frac{1}{|\Omega|}\int_{\Omega}u_{0}$. Moreover, we fix the family of approximate solutions $(u_{\varepsilon},v_{\varepsilon})_{\varepsilon\in(0,1)}$ with uniform existence time $T_{0}>0$ constructed in Lemma 2.3 as well as the limit solution $(u,v)$ found in Lemma 2.4.
@@ 3 Mass accumulation functions
on $0,R^{n}]\times[0,T_{0}]$ with initial data $w_{0}(s)\coloneqq n\int_{0}^{s^{\frac{1}{n}}}\rho^{n-1}u_{0}(\rho)\,\mathrm{d}\rho$. Additionally, it is immediately obvious from the nonnegativity of $u_{\varepsilon}$ as well as the mass conservation properties of (2.2) that $w_{\varepsilon}(\cdot,t)$ is monotonically increasing and $w_{\varepsilon}(0,t)=0$ as well as $w_{\varepsilon}(R^{n},t)=\mu R^{n}$ for all $t\in[0,T_{0}]$.
@@ 4 Comparison argument
$(\mathcal{P}_{\varepsilon}\underline{w})(s,t)$ $\leq 0$ $\qquad\text{for a.e.\ $s\in(a,b)$ and all $t\in(0,T)$},$ (4.5)
$(\mathcal{P}_{\varepsilon}\overline{w})(s,t)$ $\geq 0$ $\qquad\text{for a.e.\ $s\in(a,b)$ and all $t\in(0,T)$},$ (4.6)
$\underline{w}(s,0)-\overline{w}(s,0)$ $\leq 0$ $\qquad\text{for all $s\in(a,b)$},$ (4.7)
$\underline{w}(s,t)-\overline{w}(s,t)$ $\leq 0$ $\qquad\text{for all $s\in\{a,b\}$ and $t\in(0,T)$},$ (4.8)
then
$\underline{w}(s,t)-\overline{w}(s,t)$ $\leq 0$ $\qquad\text{for all $s\in(a,b)$ and $t\in(0,T)$}.$ (4.9)
@@
for all $(s,t)\in[0,L]\times[0,T)$ with $L\coloneqq b$. As we have only modified the original functions $\overline{w}$, $\underline{w}$ in a linear fashion, the new functions $\overline{W}$, $\underline{W}$ naturally retain all of our assumed regularity properties as well as the proper ordering at the parabolic boundary of $0,L]\times[0,T)$ due to (4.7) as well as (4.8). Due to (4.3) and $a<b$, it is further directly evident that $\partial_{s}\overline{W}>0$ and $\partial_{s}\underline{W}>0$ in $(0,L)\times(0,T)$. Finally due to (4.5) and (4.6), it follows by direct computation that
@@ 4.1 Case of support shrinking
for all $r_{0}\in[r_{\mathrm{min}},r_{1})$, and such that the quantities $\eta$ and $\delta$ defined in (4.10) fulfil
@@
for all $t\in[0,\infty)$. Moreover, $\partial_{ss}\underline{w}_{\mathrm{mid}}$ is continuous on the compact set $r_{0}^{n},\rho(t)-\delta]$ for all $t\in[0,\infty)$, so that all our desired regularity properties for $\underline{w}$ are evidently fulfilled.
@@
$\underline{w}_{\mathrm{mid}}+\mu s\leq-\theta_{\mathrm{max}}-\varepsilon_{0}\kappa(R^{n}-r_{1}^{n})-(1-\lambda)\mu(R^{n}-r_{1}^{n})$ (4.20)
@@
For our first comparison function, let now $\underline{w}_{\varepsilon,\mathrm{stat}}$ be as in (4.11) for all $\varepsilon\in(0,\varepsilon_{1})$ with the parameters as fixed above but $\theta$ replaced by 0. Then we gain
@@
Again due to $\eta\geq 0$, $\lambda>1$ as well as the fact that $w_{0}(s)\geq\mu R^{n}-C(r_{1}^{n}-s)_{+}^{\frac{m}{m-1}}\eqqcolon z_{0}(s)$ for $s\in[r_{0}^{n},R^{n}]$ by (4.21), we see that $w_{0}(s)\geq\underline{w}_{\varepsilon,\mathrm{stat}}(s,0)$ for all $s\in[r_{0}^{n},r_{1}^{n}-\delta]\cup[r_{1}^{n},R^{n}]$. Moreover, since $w_{0}\geq z_{0}$ and $z_{0}(r_{1}^{n})\geq\underline{w}_{\varepsilon,\mathrm{stat}}(r_{1}^{n},0)$, the ordering $w_{0}(s)\geq\underline{w}_{\varepsilon,\mathrm{stat}}(s,0)$ in $r_{1}^{n}-\delta,r_{1}^{n}]$ and hence on the full interval $r_{0}^{n},R^{n}]$ follows if $\underline{w}_{\varepsilon,\mathrm{stat}}(\cdot,0)-z_{0}$ is increasing in the former interval. Indeed,
@@
Notably when $\theta=0$, the condition (4.12) is always trivially fulfilled. Thus Lemma 4.2 ensures that we can apply the comparison result in Lemma 4.1 to $w_{\varepsilon}$ and $\underline{w}_{\varepsilon,\mathrm{stat}}$ and thus gain $w_{\varepsilon}\geq\underline{w}_{\varepsilon,\mathrm{stat}}$ on $r_{0}^{n},R^{n}]\times[0,T]$. This then directly yields
@@
a.e. on $r_{0}^{n},r_{2}^{n}]\times[0,T]$ due to Lemma 4.2 and our previous parameter choices. As initial data ordering follows by the same argument as before, this again allows us to apply Lemma 4.1 to gain that
@@
on $r_{0}^{n},r_{1}^{n}]\times[0,T]$. As $\underline{w}_{\varepsilon,\mathrm{shrink}}$ converges pointwise to